OpenStudy (anonymous):

y = x + 4 –3x + 2y = 8 1) identify the slope and the y-intercept of the equation. 3.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 4.Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. 5.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 6.In a paragraph of at least three sentences, explain how the graphs of the two equequations in your group are the same and how they are different. Be sure to use complete sentences

OpenStudy (anonymous):

put in slope intercept form y= mx+b b= y intercept y = x + 4 already in slope intercept form –3x + 2y = 8 solve for y in second equation –3x + 2y = 8 1/2 [ 2y = 3x+ 8 ] y = (3/2)x + 4

OpenStudy (anonymous):

thanks for your help, but what about question 5, and 6

OpenStudy (anonymous):

well what can you say about those equations?

OpenStudy (anonymous):

sorry the lines

OpenStudy (anonymous):

what directions are they going up or down

OpenStudy (anonymous):



OpenStudy (anonymous):

this is not to scale

OpenStudy (anonymous):

then i think you can describe this graph in your own words right